grep, searching for a pattern

version 2.3, 10 February 1999

Alain Magloire et al.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1998 Free Software Foundation, Inc.

Published by the Free Software Foundation, 59 Temple Place - Suite 330, Boston, MA 02111-1307, USA Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Foundation.

1 Introduction

grep searches the input files for lines containing a match to a given pattern list. When it finds a match in a line, it copies the line to standard output (by default), or does whatever other sort of output you have requested with options. grep expects to do the matching on text. Since newline is also a separator for the list of patterns, there is no way to match newline characters in a text.

2 Invoking grep

grep comes with a rich set of options from POSIX.2 and GNU extensions.

‘-c’

‘--count’ Suppress normal output; instead print a count of matching lines for each input file. With the ‘-v’, ‘--revert-match’ option, count non-matching lines.

‘-e pattern’

‘--regexp=pattern’

Use pattern as the pattern; useful to protect patterns beginning with a ‘-’.

‘-f file’

‘--file=file’

Obtain patterns from file, one per line. The empty file contains zero patterns, and therefore matches nothing.

‘-i’
‘--ignore-case’

Ignore case distinctions in both the pattern and the input files.

‘-l’
‘--files-with-matches’

Suppress normal output; instead print the name of each input file from which output would normally have been printed. The scanning of every file will stop on the first match.

‘-n’
‘--line-number’

Prefix each line of output with the line number within its input file.

‘-q’
‘--quiet’
‘--silent’

Quiet; suppress normal output. The scanning of every file will stop on the first match. Also see the ‘-s’ or ‘--no-messages’ option.

‘-s’
‘--no-messages’

Suppress error messages about nonexistent or unreadable files. Portability note: unlike GNU grep, BSD grep does not comply with POSIX.2, because BSD grep lacks a ‘-q’ option and its ‘-s’ option behaves like GNU grep’s ‘-q’ option. Shell scripts intended to be portable to BSD grep should avoid both ‘-q’ and ‘-s’ and should redirect output to ‘/dev/null’ instead.

‘-v’
‘--revert-match’

Invert the sense of matching, to select non-matching lines.

‘-x’
‘--line-regexp’

Select only those matches that exactly match the whole line.

2.1 GNU Extensions

‘-A num’

‘--after-context=num’

Print num lines of trailing context after matching lines.

‘-B num’

‘--before-context=num’

Print num lines of leading context before matching lines.

‘-C’
‘--context[=num]’

Print num lines (default 2) of output context.

‘-NUM’ Same as ‘--context=num’ lines of leading and trailing context. However, grep will never print any given line more than once.

‘-V’
‘--version’

Print the version number of grep to the standard output stream. This version number should be included in all bug reports.

‘--help’ Print a usage message briefly summarizing these command-line options and the bug-reporting address, then exit.

‘-b’
‘--byte-offset’

Print the byte offset within the input file before each line of output. When grep runs on MS-DOS or MS-Windows, the printed byte offsets depend on whether the ‘-u’ (‘--unix-byte-offsets’) option is used; see below.

‘-d action’

‘--directories=action’

If an input file is a directory, use action to process it. By default, action is ‘read’, which means that directories are read just as if they were ordinary files (some operating systems and filesystems disallow this, and will cause grep to print error messages for every directory). If action is ‘skip’, directories are silently skipped. If action is ‘recurse’, grep reads all files under each directory, recursively; this is equivalent to the ‘-r’ option.

‘-h’
‘--no-filename’

Suppress the prefixing of filenames on output when multiple files are searched.

‘-L’
‘--files-without-match’

Suppress normal output; instead print the name of each input file from which no output would normally have been printed. The scanning of every file will stop on the first match.

‘-a’

‘--text’ Do not suppress output lines that contain binary data. Normally, if the first few bytes of a file indicate that the file contains binary data, grep outputs only a message saying that the file matches the pattern. This option causes grep to act as if the file is a text file, even if it would otherwise be treated as binary. Warning: the result might be binary garbage printed to the terminal, which can have nasty side-effects if the terminal driver interprets some of it as commands.

‘-w’
‘--word-regexp’

Select only those lines containing matches that form whole words. The test is that the matching substring must either be at the beginning of the line, or preceded by a non-word constituent character. Similarly, it must be either at the end of the line or followed by a non-word constituent character. Wordconstituent characters are letters, digits, and the underscore.

‘-r’
‘--recursive’

For each directory mentioned in the command line, read and process all files in that directory, recursively. This is the same as the ‘-d recurse’ option.

‘-y’ Obsolete synonym for ‘-i’.

‘-U’
‘--binary’

Treat the file(s) as binary. By default, under MS-DOS and MS-Windows, grep guesses the file type by looking at the contents of the first 32KB read from the file. If grep decides the file is a text file, it strips the CR characters from the original file contents (to make regular expressions with ^and $ work correctly). Specifying ‘-U’ overrules this guesswork, causing all files to be read and passed to the matching mechanism verbatim; if the file is a text file with CR/LF pairs at the end of each line, this will cause some regular expressions to fail. This option is only supported on MS-DOS and MS-Windows.

‘-u’
‘--unix-byte-offsets’

Report Unix-style byte offsets. This switch causes grep to report byte offsets as if the file were Unix style text file, i.e. the byte offsets ignore the CR characters which were stripped off. This will produce results identical to running grep on a Unix machine. This option has no effect unless ‘-b’ option is also used; it is only supported on MS-DOS and MS-Windows.

Several additional options control which variant of the grep matching engine is used. See Chapter 4 [Grep Programs], page 4.

grep uses the environment variable LANG to provide internationalization support, if compiled with this feature.

3 Diagnostics

Normally, exit status is 0 if matches were found, and 1 if no matches were found (the ‘-v’ option inverts the sense of the exit status). Exit status is 2 if there were syntax errors in the pattern, inaccessible input files, or other system errors.

4 grep programs

grep searches the named input files (or standard input if no files are named, or the file name ‘-’ is given) for lines containing a match to the given pattern. By default, grep prints the matching lines. There are three major variants of grep, controlled by the following options.

‘-G’
‘--basic-regexp’

Interpret pattern as a basic regular expression. This is the default.

‘-E’
‘--extended-regexp’

Interpret pattern as an extended regular expression.

‘-F’
‘--fixed-strings’

Interpret pattern as a list of fixed strings, separated by newlines, any of which is to be matched.

In addition, two variant programs egrep and fgrep are available. egrep is similar (but not identical) to ‘grep -E’, and is compatible with the historical Unix egrep. fgrep is the same as ‘grep -F’.

5 Regular Expressions

A regular expression is a pattern that describes a set of strings. Regular expressions are constructed analogously to arithmetic expressions, by using various operators to combine smaller expressions. grep understands two different versions of regular expression syntax: “basic” and “extended”. In GNU grep, there is no difference in available functionality using either syntax. In other implementations, basic regular expressions are less powerful. The following description applies to extended regular expressions; differences for basic regular expressions are summarized afterwards.

The fundamental building blocks are the regular expressions that match a single character. Most characters, including all letters and digits, are regular expressions that match themselves. Any metacharacter with special meaning may be quoted by preceding it with a backslash. A list of characters enclosed by ‘[’ and ‘]’ matches any single character in that list; if the first character of the list is the caret ‘^’, then it matches any character not in the list. For example, the regular expression ‘[0123456789]’ matches any single digit. A range of ascii characters may be specified by giving the first and last characters, separated by a hyphen. Finally, certain named classes of characters are predefined. Their names are self explanatory, and they are :

‘[:alnum:]’

Any of [:digit:] or [:alpha:]

‘[:alpha:]’

Any local-specific or one of the ascii letters:

          a b c d e f g h i j k l m n o p q r s t u v w x y z,
          A B C D E F G H I J K L M N O P Q R S T U V W X Y Z.
‘[:cntrl:]’
          Any of BEL, BS, CR, FF, HT, NL, or VT.
‘[:digit:]’
          Any one of 0 1 2 3 4 5 6 7 8 9.
‘[:graph:]’

Anything that is not a ‘[:alphanum:]’ or ‘[:punct:]’.

‘[:lower:]’
          Any one of a b c d e f g h i j k l m n o p q r s t u v w x y z.
‘[:print:]’

Any character from the ‘[:space:]’ class, and any character that is not in the

          ‘[:isgraph:]’ class.
‘[:punct:]’
          Any one of ! " #% & ' ( ) ; < = > ? [ \ ] * + , - . / : ^_{ | }.
‘[:space:]’
          Any one of CR FF HT NL VT SPACE.
‘[:upper:]’
          Any one of A B C D E F G H I J K L M N O P Q R S T U V W X Y Z.
‘[:xdigit:]’
          Any one of a b c d e f A B C D E F 0 1 2 3 4 5 6 7 8 9.

For example, ‘[[:alnum:]]’ means ‘[0-9A-Za-z]’, except the latter form is dependent upon the ascii character encoding, whereas the former is portable. (Note that the brackets in these class names are part of the symbolic names, and must be included in addition to the brackets delimiting the bracket list). Most metacharacters lose their special meaning inside lists. To include a literal ‘]’, place it first in the list. Similarly, to include a literal ‘^’, place it anywhere but first. Finally, to include a literal ‘-’, place it last.

The period ‘.’ matches any single character. The symbol ‘\w’ is a synonym for

‘[[:alnum:]]’ and ‘\W’ is a synonym for ‘[^[:alnum]]’.

The caret ‘^’ and the dollar sign ‘$’ are metacharacters that respectively match the empty string at the beginning and end of a line. The symbols ‘\<’ and ‘\>’ respectively match the empty string at the beginning and end of a word. The symbol ‘\b’ matches the empty string at the edge of a word, and ‘\B’ matches the empty string provided it’s not at the edge of a word.

A regular expression may be followed by one of several repetition operators:

‘?’ The preceding item is optional and will be matched at most once.

‘*’ The preceding item will be matched zero or more times.

‘+’ The preceding item will be matched one or more times.

‘{n}’ The preceding item is matched exactly n times.

‘{n,}’ The preceding item is matched n or more times.

‘{,m}’ The preceding item is optional and is matched at most m times.

‘{n,m}’ The preceding item is matched at least n times, but not more than m times.

Two regular expressions may be concatenated; the resulting regular expression matches any string formed by concatenating two substrings that respectively match the concatenated subexpressions.

Two regular expressions may be joined by the infix operator ‘|’; the resulting regular expression matches any string matching either subexpression.

Repetition takes precedence over concatenation, which in turn takes precedence over alternation. A whole subexpression may be enclosed in parentheses to override these precedence rules.

The backreference ‘\n’, where n is a single digit, matches the substring previously matched by the nth parenthesized subexpression of the regular expression.

In basic regular expressions the metacharacters ‘?’, ‘+’, ‘{’, ‘|’, ‘(’, and ‘)’ lose their special meaning; instead use the backslashed versions ‘\?’, ‘\+’, ‘\{’, ‘\|’, ‘\(’, and ‘\)’.

In egrep the metacharacter ‘{’ loses its special meaning; instead use ‘\{’. This not true

for ‘grep -E’.

6 Reporting bugs

Email bug reports to bug-gnu-utils@gnu.org. Be sure to include the word “grep” somewhere in the “Subject:” field.

Large repetition counts in the ‘{m,n}’ construct may cause grep to use lots of memory. In addition, certain other obscure regular expressions require exponential time and space, and may cause grep to run out of memory. Backreferences are very slow, and may require exponential time.

Concept Index 7

Concept Index

This is a general index of all issues discussed in this manual, with the exception of the grep commands and command-line options.

A

after context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 alphabetic characters. . . . . . . . . . . . . . . . . . . . . . . . . . 5 alphanumeric characters . . . . . . . . . . . . . . . . . . . . . . . 5 asterisk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

B

basic regular expressions . . . . . . . . . . . . . . . . . . . . . . 6 before context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 binary files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 binary files, DOS/Windows . . . . . . . . . . . . . . . . . . . . 3 braces, first argument omitted . . . . . . . . . . . . . . . . . 6 braces, one argument. . . . . . . . . . . . . . . . . . . . . . . . . . 6 braces, second argument omitted . . . . . . . . . . . . . . . 6 braces, two arguments. . . . . . . . . . . . . . . . . . . . . . . . . 6 Bugs, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 byte offset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 byte offsets, on DOS/Windows. . . . . . . . . . . . . . . . . 3

C

case insensitive search . . . . . . . . . . . . . . . . . . . . . . . . . 1 case insensitive search, obsolete option. . . . . . . . . . 3 character classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 classes of characters . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 context lines, after match. . . . . . . . . . . . . . . . . . . . . . 2 context lines, before match . . . . . . . . . . . . . . . . . . . . 2 control characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 counting lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

D

digit characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 directory search. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 DOS byte offsets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 DOS/Windows binary files . . . . . . . . . . . . . . . . . . . . 3

F

files which don’t match . . . . . . . . . . . . . . . . . . . . . . . . 3

G

graphic characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

H

hexadecimal digits . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

L

line numbering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 lower-case alphabetic characters. . . . . . . . . . . . . . . . 5

M

match sub-expression at most m times. . . . . . . . . . 6 match sub-expression at most once . . . . . . . . . . . . . 6 match sub-expression n or more times . . . . . . . . . . 6 match sub-expression n times . . . . . . . . . . . . . . . . . . 6 match sub-expression zero or more times . . . . . . . 6 match the whole line . . . . . . . . . . . . . . . . . . . . . . . . . . 2 matching basic regular expressions . . . . . . . . . . . . . 4 matching extended regular expressions. . . . . . . . . . 4 matching fixed strings . . . . . . . . . . . . . . . . . . . . . . . . . 4 matching whole words. . . . . . . . . . . . . . . . . . . . . . . . . 3

N

names of matching files. . . . . . . . . . . . . . . . . . . . . . . . 1 no filename prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 numeric characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

P

pattern from file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 pattern list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 plus sign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 print non-matching lines. . . . . . . . . . . . . . . . . . . . . . . 2 printable characters . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 punctuation characters . . . . . . . . . . . . . . . . . . . . . . . . 5

Q

question mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 quiet, silent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Concept Index 8

R

recursive search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 regular expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 revert matching. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

S

searching directory trees. . . . . . . . . . . . . . . . . . . . . . . 3 Searching for a pattern. . . . . . . . . . . . . . . . . . . . . . . . 1 space characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 suppress binary data . . . . . . . . . . . . . . . . . . . . . . . . . . 3 suppress error messages . . . . . . . . . . . . . . . . . . . . . . . 1

U

upper-case alphabetic characters . . . . . . . . . . . . . . . 5 Usage summary, printing . . . . . . . . . . . . . . . . . . . . . . 2

V

Version, printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

W

whitespace characters . . . . . . . . . . . . . . . . . . . . . . . . . 5

X

xdigit class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Index 9

Index

This is an alphabetical list of all grep commands and command-line options.

*

* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

-

--after-context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --basic-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 --before-context . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --binary. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 --byte-offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --context. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --extended-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 --file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 --files-with-matches. . . . . . . . . . . . . . . . . . . . . . . . 1 --files-without-match. . . . . . . . . . . . . . . . . . . . . . . 3 --fixed-strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 --help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --ignore-case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 --line-number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 --line-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --no-filename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --no-messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 --quiet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 --recursive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 --regexp=pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 --revert-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --silent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 --text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 --unix-byte-offsets . . . . . . . . . . . . . . . . . . . . . . . . . 3 --version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --word-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 -C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 -d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 -E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 -f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 -F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 -G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 -h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 -l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 -L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 -NUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 -r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 -u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -U . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

?

? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

                              {

{,m} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 {n,} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 {n,m} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 {n} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

                              +

+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

A

alnum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 alpha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

C

cntrl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

D

digit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Index 10

G

graph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

L

lower . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

P

print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 punct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

S

space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

U

upper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

X

xdigit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Invoking grep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2.1 GNU Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Diagnostics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4 grep programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 5 Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . 4 6 Reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9